OpenStudy (anonymous):

Let f(x) = 3x – 42 and g(x) = 5x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

If this is a test, you should go ahead and do these problems on your own.

OpenStudy (anonymous):

It's not a test

OpenStudy (anonymous):

mhmm

OpenStudy (zepp):

You can find the answer when you cover x with some fog.

OpenStudy (anonymous):

I need help. help me through each step please idk where to start

OpenStudy (zepp):

\(\large (f ◦ g)(x)=f(g(x))\)

OpenStudy (zepp):

If I have \(\large f(x)=2x;~~g(x)=14x+8\) then \(\large f(g(x))=2(14x+8)\) That's the maximum hint I can give.

OpenStudy (anonymous):

Katie, \( f\circ g\) is read as "f of g of x" and any time you say "of" it means you're plugging something in. In this case, g(x) gets plugged into f(x). So look at your f(x) function, anywhere you see an x, plug in the whole g(x) function with some parenthesis around it.

OpenStudy (anonymous):

The equation is the same. The argument can change. if f(x) = 3*x - 42, and you want to get f(74*y^3), then you would replace x with 74*y^3 to get: f(74*y^3) = 3*(74*y^3) - 42. All you do is replace the original argument with whatever you want it to be. What we did was replace x with g(x) to get f(g(x)) = 3 * g(x) - 42. Since we knew that g(x) was equal to 5*x, we then went in the equation and replaced g(x) with 5*x to get f(g(x)) = 3 * (5 * x) - 42. We then simplified to get f(g(x)) = 15 * x - 42.

OpenStudy (anonymous):

hop that help

OpenStudy (anonymous):

thanks so much